# Composing Distributed Computations Through Task and Kernel Fusion

### v1

@@ Abstract.
We introduce Diffuse, a system that dynamically performs task and kernel fusion in distributed, task-based runtime systems. The key component of Diffuse is an intermediate representation of distributed computation that enables the necessary analyses for the fusion of distributed tasks to be performed in a scalable manner. We pair task fusion with a JIT compiler to fuse together the kernels within fused tasks. We show empirically that Diffuse’s intermediate representation is general enough to be a target for two real-world, task-based libraries (cuNumeric and Legate Sparse), letting Diffuse find optimization opportunities across function and library boundaries. Diffuse accelerates unmodified applications developed by composing task-based libraries by 1.86x on average (geo-mean), and by between 0.93x–10.7x on up to 128 GPUs. Diffuse also finds optimization opportunities missed by the original application developers, enabling high-level Python programs to match or exceed the performance of an explicitly parallel MPI library.
@@ 1. Introduction
However, the same abstractions that yield important composition properties internally and externally to these distributed libraries can result in degraded end-to-end performance. The task decomposition of library operations results in tasks that may be optimized individually but can have poor data locality and allocate much more temporary data than a different program organization that breaks down the abstraction boundaries by fusing tasks together both within the operations of a particular library and across library boundaries. As the task-based runtime system is issued a stream of tasks after library abstraction boundaries have been traversed, the runtime has the opportunity to fuse the tasks from different libraries together, which in turn enables the fusion of the kernels of nested loops within fused tasks. Fusion by the task-based runtime allows for these optimizations to be performed without being limited to the semantics of any particular library.
@@
We present Diffuse, a system that performs task and kernel fusion for distributed, task-based runtime systems, transparently achieving optimizations found in hand-tuned programs. Diffuse reasons over a task-based intermediate representation (IR) of distributed computation, modeling computation as a sequence of tasks operating on partitioned data (Section 3). Diffuse’s IR is scale-free, meaning that the size of the IR and analyses on it are independent of the size of the target machine. Diffuse uses this IR to dynamically fuse tasks in a distributed-memory setting (Section 4). Diffuse pairs task fusion with a JIT compiler based on MLIR (Lattner et al., 2020) that fuses and optimizes kernels within fused tasks, enabling data reuse across independent tasks (Section 6). By analyzing a task-based IR, Diffuse’s optimizations are not tied to the semantics of any particular library.
We implement Diffuse as a middle layer between high-level task-based libraries and the low-level Legion runtime system (Bauer et al., 2012). To demonstrate Diffuse, we modify the implementations of the distributed libraries cuNumeric (Bauer and Garland, 2019) and Legate Sparse (Yadav et al., 2023) to target Diffuse’s IR, and to expose their task implementations in MLIR for Diffuse’s compiler to process. Diffuse then performs dynamic analyses to fuse the tasks and kernels issued by these libraries before forwarding the optimized tasks to Legion. As a result, programmers using cuNumeric and Legate Sparse benefit from Diffuse without modifying their applications.
To evaluate Diffuse, we apply it to micro-benchmarks and several full scientific computing applications developed in cuNumeric and Legate Sparse, including sparse Krylov solvers and physical simulations. We compare against the standard implementations of cuNumeric and Legate Sparse and show that Diffuse achieves 1.86x speedup on average (geo-mean) over unmodified applications on up to 128 GPUs. We additionally compare against the high-performance MPI-based PETSc (Balay et al., 2022) library and show that Diffuse enables naturally written NumPy and SciPy Sparse programs to match or exceed the performance of PETSc (1.4x geo-mean speedup). Finally, we show that Diffuse is able to find fusion and optimization opportunities missed by the original application developers, achieving 1.23x speedup on average (geo-mean) over already hand-optimized code.
@@ 2. Motivating Example
(a) cuNumeric source code.
@@
Figure 2. Execution example of Diffuse on a distributed, multi-GPU cuNumeric 5-point stencil application.
@@
Figure 2 shows how Diffuse optimizes the cuNumeric program in Figure 2(a) that performs a 5-point stencil computation. The cuNumeric library is a drop-in replacement for NumPy (Harris et al., 2020) that scales unmodified NumPy programs to distributed machines by targeting the Legion (Bauer et al., 2012) runtime system. cuNumeric maps NumPy arrays to Legion’s regions, and maps NumPy functions to task launches operating on regions that are partitioned across the machine. The program execution on a four-by-four grid with four nodes is visualized in Figure 2(b), where each node owns an element of each aliasing view of the grid array. The dotted arrows represent the communication required to propagate updates to the center array to the other aliasing views of grid. Figure 2(c) is a simplified representation of the task stream that cuNumeric issues during execution of the inner loop (lines 1014 of Figure 2(a)), and Figure 2(e) contains pseudocode for each of the task implementations. This stream of operations creates multiple temporary distributed arrays for the results of individual operations, and separate tasks for each corresponding addition and multiplication. The combination of temporary arrays and separate tasks of loops is an inefficient execution strategy. Diffuse speeds this program up by four times by creating a new fused task that computes the work array (lines 11-13) in a single operation and removes the temporary arrays, including avg, resulting in the stream of operations in Figure 2(d) and the generated fused task in Figure 2(f). Interestingly, Diffuse does not fuse the task that performs center[:] = work (line 14 of Figure 2(a)).
@@
To understand these decisions, we must introduce the distributed aspect of the tasks and data collections in Figure 2(c). Each task in Figure 2(c) actually represents a group of parallel tasks launched over partitioned arrays, where each parallel task operates on a subset of the partitioned data. Dependencies and communication that arise from parallel tasks operating on the same distributed data affect when fusion is possible. In our example, the arrays center, north, east, west, and south are aliasing views of the array grid, meaning that they share logical array entries. Because these distributed arrays alias, Diffuse does not fuse the task group that computes center[:] = work into the task group that reads from north, east, west and south, as the fusion would create a task group that concurrently reads and writes to aliasing data. Similarly, the center[:] = work task group issued at iteration $i$ cannot be fused into the avg computation (line 11 of Figure 2(a)) at iteration $i+1$ because communication is required to propagate updates to center. To reason about distributed computations over partitioned data, we develop a scale-free intermediate representation (Section 3) that models tasking runtime systems which support aliased views of distributed data. We then develop an algorithm for task fusion (Section 4) that reasons about communication in distributed computations to fuse groups of parallel tasks.
@@ 3. Intermediate Representation
The first contribution of Diffuse is an IR that enables scalable fusion analyses through a scale-free representation of distributed programs, meaning that the size of the representation is independent of the total number of processors in the target system. Diffuse’s IR is an abstraction over the collections of concrete tasks and distributed data structures of a lower-level task-based programming system, like Legion, that usually have scale-aware representations. We have modified cuNumeric and Legate Sparse to target Diffuse’s IR instead of targeting Legion directly. Diffuse’s IR, presented in Figure 3, is designed to make it inexpensive to perform the analyses required for fusion, while still being able to express sophisticated computations. The IR contains a data model to represent distributed data, and a computational model to define distributed computations over distributed data. The syntax of the IR is in Figure 3(a), and a visualization of the IR’s structure is shown in Figure 3(b).
3.1. Data Model
@@ 3.2. Computational Model
Diffuse models computation as a stream of index tasks (Soi et al., 2021) issued in program order. An $\textsf{IndexTask}(d,A)$ represents a group of parallel tasks over points in a rectangular launch domain $d$. An index task operates on the list $A$ of stores, partitions, and privileges, using the denoted privilege to access the requested partition of each store. We refer to each privilege with the abbreviations noted in parentheses. Each parallel task within the group reads from, writes to, or reduces to the sub-stores referred to by the stores and partitions at each point. For the simplicity of presentation, we assume that the Reduce privilege refers to a single reduction function being applied (such as addition). This representation is explicitly parallel as tasks are annotated with their launch domain and partitions of distributed data structures. However, the representation is scale-free as the size of the representation is independent of the degree of parallelism — only the symbolic size of the launch domain increases.
@@ 4.3. Proof of Correctness
Given a window of tasks $[T_{1},\ldots,T_{n}]$, our task fusion algorithm identifies a prefix $[T_{1},\ldots,T_{f}]$ and produces a fused task $F$ such that $[T_{1},\ldots,T_{f}]$ are fusible and $F$ preserves all dependencies of the task sequence $[T_{1},\ldots,T_{f}]$.
Proof Sketch. To show that $[T_{1},\ldots,T_{f}]$ are fusible we proceed by contradiction and show that if any dependencies in Definition 1 exist, then a fusion constraint must have been violated. Since $[T_{1},\ldots,T_{f}]$ are fusible, the dependence between any $T_{i}^{p}$ and $T_{j}^{p}$ must be pointwise, so program order fusion of $[T_{1},\ldots,T_{f}]$ preserves all dependencies.
@@ 5.1. Temporary Store Elimination
(a) cuNumeric code fragment.
@@
To introduce when a store is temporary, consider the cuNumeric program in Figure 8(a) and the resulting task stream in Figure 8(b). This example introduces some new operations, specifically flush_window, which sends all pending tasks through Diffuse to the underlying runtime system, and the Python del operator, which drops references. The program creates the stores x, y, z, w, and v. Consider the program state after line 10: the tasks that initialize x and y have executed, as the first flush_window call sent those tasks to Diffuse. The fusion algorithm determines that the tasks issued by lines 46 can be fused, while the final norm must be excluded. First, v is not temporary because the application holds a reference to it. Next, while there are no application references to w, the pending norm task reads a portion of w, making it not temporary. The stores x and y are only read by the fused task, and thus are not temporary. Only z is temporary because it is produced entirely within the fused task and is not visible to the application or pending tasks. We formalize this intuition as constraints that must be satisfied for a store to be temporary.
@@
The function $\textsf{covers}(S,P)$ is true when the partition $P$ contains all points in the store $S$. The first two constraints check that the store’s contents are entirely created within the fused task and not used by any other existing task. The third constraint ensures that the application can no longer view any effects on a store, checked through a split reference counting scheme in the implementation of Diffuse’s IR. Temporary stores are demoted from a distributed allocation into a task-local allocation, as described in Section 6.
@@ 6. Kernel Fusion
The final component of Diffuse is a compilation stack to optimize fused tasks. A high-level program representation is required to both perform optimizations like loop fusion and to lower to different backends like GPUs and multi-threaded CPUs. We leverage the MLIR compiler stack, which is extensible and is pre-packaged with many common compiler analyses. We first provide background on MLIR, and then describe the code generation process and optimizations performed within Diffuse.
@@ 6.2. Generator Functions
To describe Diffuse’s compiler, we walk through the stages that a fused task traverses. cuNumeric and Legate Sparse developers implement tasks by defining variants that target CPUs or GPUs. To use Diffuse, developers register a generator function with Diffuse that returns an MLIR fragment describing the task’s computation. We found the integration effort of adding these generator functions to be modest, requiring 50–100 lines of C++ code per operation. We emphasize that only library developers, not end users, must develop MLIR kernels for tasks. Additionally, the integration effort was incremental—as more tasks were implemented with MLIR generators, Diffuse could exploit more fusion. An example generated MLIR fragment by cuNumeric for an element-wise addition operation is shown in Figure 12(a).
@@ 6.3. Compilation Pipeline
When Diffuse identifies that a sequence of tasks may be fused, it invokes each task generator and constructs an MLIR module containing the body of each task in the original program order. Figure 12(b) shows a fused task for the cuNumeric computation c = a + b; e = c + d, where all variables represent distributed vectors. This program originally has two index tasks (one for each add operation) which are fused into a single index task where the original task bodies (the MLIR in Figure 12(a)) appear sequentially in the fused task. Before optimization of the task body, Diffuse first promotes distributed data into task-local allocations, resulting in Figure 12(c).

### v2

@@ Abstract.
We introduce Diffuse, a system that dynamically performs task and kernel fusion in distributed, task-based runtime systems. The key component of Diffuse is an intermediate representation of distributed computation that enables the necessary analyses for the fusion of distributed tasks to be performed in a scalable manner. We pair task fusion with a JIT compiler to fuse together the kernels within fused tasks. We show empirically that Diffuse’s intermediate representation is general enough to be a target for two real-world, task-based libraries (cuPyNumeric and Legate Sparse), letting Diffuse find optimization opportunities across function and library boundaries. Diffuse accelerates unmodified applications developed by composing task-based libraries by 1.86x on average (geo-mean), and by between 0.93x–10.7x on up to 128 GPUs. Diffuse also finds optimization opportunities missed by the original application developers, enabling high-level Python programs to match or exceed the performance of an explicitly parallel MPI library.
Distributed Programming; Composable Software
journalyear: 2025copyright: ccconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 30-April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’25), March 30-April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3669940.3707216isbn: 979-8-4007-0698-1/25/03ccs: Computing methodologies Distributed programming languages
@@ 1. Introduction
However, the same abstractions that yield important composition properties internally and externally to these distributed libraries can result in degraded end-to-end performance. The task decomposition of library operations results in tasks that may be optimized individually but can have poor data locality and allocate much more temporary data than a different program organization that breaks down the abstraction boundaries by fusing tasks together both within the operations of a particular library and across library boundaries. As the task-based runtime system is issued a stream of tasks after library abstraction boundaries have been traversed, the runtime has the opportunity to fuse the tasks from different libraries together, which in turn enables the fusion of the kernels of nested loops within fused tasks. Fusion by the task-based runtime allows for these optimizations to be performed without being limited to the semantics of any particular task-based library.
@@
We present Diffuse, a system that dynamically performs task and kernel fusion for distributed, task-based runtime systems, transparently achieving optimizations found in hand-tuned programs. Diffuse reasons over a task-based IR of distributed computation, modeling computation as a sequence of tasks operating on partitioned data (Section 3). Diffuse’s IR is scale-free, meaning that the size of the IR and analyses on it are independent of the size of the target machine. Diffuse uses this IR to perform a dynamic dependence analysis to fuse tasks in a distributed-memory setting (Section 4). Diffuse pairs task fusion with a JIT compiler based on MLIR (Lattner et al., 2020) that fuses and optimizes kernels within dynamically fused tasks, enabling data reuse across independent tasks (Section 6). By analyzing a task-based IR, Diffuse’s optimizations are not tied to the semantics of any particular library.
We implement Diffuse as a middle layer between high-level task-based libraries and the low-level Legion runtime system (Bauer et al., 2012). To demonstrate Diffuse, we modify the implementations of the distributed libraries cuPyNumeric (Bauer and Garland, 2019) and Legate Sparse (Yadav et al., 2023) to target Diffuse’s IR, and to expose their task implementations in MLIR for Diffuse’s compiler to process. Diffuse then performs dynamic analyses to fuse the tasks and kernels issued by these libraries before forwarding the optimized tasks to Legion. As a result, programmers using cuPyNumeric and Legate Sparse benefit from Diffuse without modifying their applications.
To evaluate Diffuse, we apply it to micro-benchmarks and several full scientific computing applications developed in cuPyNumeric and Legate Sparse, including sparse Krylov solvers and physical simulations. We compare against the standard implementations of cuPyNumeric and Legate Sparse and show that Diffuse achieves 1.86x speedup on average (geo-mean) over unmodified applications on up to 128 GPUs. We additionally compare against the high-performance MPI-based PETSc (Balay et al., 2022) library and show that Diffuse enables naturally written NumPy and SciPy Sparse programs to match or exceed the performance of PETSc (1.4x geo-mean speedup). Finally, we show that Diffuse is able to find fusion and optimization opportunities missed by the original application developers, achieving 1.23x speedup on average (geo-mean) over already hand-optimized code.
@@ 2. Motivating Example
(a) cuPyNumeric source code.
@@
Figure 2. Execution example of Diffuse on a distributed, multi-GPU cuPyNumeric 5-point stencil application.
@@
Figure 2 shows how Diffuse optimizes the cuPyNumeric program in Figure 2(a) that performs a 5-point stencil computation. The cuPyNumeric library is a drop-in replacement for NumPy (Harris et al., 2020) that scales unmodified NumPy programs to distributed machines by targeting the Legion (Bauer et al., 2012) runtime system. cuPyNumeric maps NumPy arrays to Legion’s regions, and maps NumPy functions to task launches operating on regions that are partitioned across the machine. As the cuPyNumeric program executes, it issues a stream of tasks to the Legion runtime, which dynamically discovers the necessary communication and synchronization required to execute the tasks on the target machine. The program execution on a four-by-four grid with four nodes is visualized in Figure 2(b), where each node owns an element of each aliasing view of the grid array. The dotted arrows represent the communication required to propagate updates to the center array to the other aliasing views of grid. Figure 2(c) is a simplified representation of the task stream that cuPyNumeric issues during execution of the inner loop (lines 1014 of Figure 2(a)), and Figure 2(e) contains pseudocode for each of the task implementations. This stream of operations creates multiple temporary distributed arrays for the results of individual operations, and separate tasks for each corresponding addition and multiplication. The combination of temporary arrays and separate tasks of loops is an inefficient execution strategy. Diffuse speeds this program up by four times by creating a new fused task that computes the work array (lines 11-13) in a single operation and removes the temporary arrays, including avg, resulting in the stream of operations in Figure 2(d) and the generated fused task in Figure 2(f). Interestingly, Diffuse does not fuse the task that performs center[:] = work (line 14 of Figure 2(a)).
@@
To understand these decisions, we must introduce the distributed aspect of the tasks and data collections in Figure 2(c). Each task in Figure 2(c) actually represents a group of parallel tasks launched over partitioned arrays, where each parallel task operates on a subset of the partitioned data. Dependencies and communication that arise from parallel tasks operating on the same distributed data affect when fusion is possible. In our example, the arrays center, north, east, west, and south are aliasing views of the array grid, meaning that they share logical array entries. Because these distributed arrays alias, Diffuse does not fuse the task group that computes center[:] = work into the task group that reads from north, east, west and south, as the fusion would create a task group that concurrently reads and writes to aliasing data. Similarly, the center[:] = work task group issued at iteration $i$ cannot be fused into the avg computation (line 11 of Figure 2(a)) at iteration $i+1$ because communication is required to propagate updates to center. To reason about distributed computations over partitioned data, we develop a scale-free intermediate representation (Section 3) that models tasking runtime systems which support aliased views of distributed data. We then develop a dynamic analysis for task fusion (Section 4) that reasons about dynamically known communication patterns in distributed computations to fuse groups of parallel tasks.
@@ 3. Intermediate Representation
The first contribution of Diffuse is an IR that enables scalable fusion analyses through a scale-free representation of distributed programs, meaning that the size of the representation is independent of the total number of processors in the target system. Diffuse’s IR is an abstraction over the collections of concrete tasks and distributed data structures of a lower-level task-based programming system, like Legion, that usually have scale-aware representations. We have modified cuPyNumeric and Legate Sparse to dynamically generate programs in Diffuse’s IR instead of targeting Legion directly. Diffuse’s IR, presented in Figure 3, is designed to make it inexpensive to perform the analyses required for fusion, while still being able to express sophisticated computations. The IR contains a data model to represent distributed data, and a computational model to define distributed computations over distributed data. The syntax of the IR is in Figure 3(a), and a visualization of the IR’s structure is shown in Figure 3(b).
3.1. Data Model
@@ 3.2. Computational Model
Diffuse models computation as a stream of index tasks (Soi et al., 2021) issued in program order. An $\textsf{IndexTask}(d,A)$ represents a group of parallel tasks over points in a rectangular domain $d$, referred to as the launch domain. An index task operates on the list $A$ of stores, partitions, and privileges, using the denoted privilege to access the requested partition of each store. We refer to each privilege with the abbreviations noted in parentheses. Each parallel task within the group reads from, writes to, or reduces to the sub-stores referred to by the stores and partitions at each point. The parallel tasks within an index task group may perform arbitrary computation on argument stores that respects the requested privilege on each argument store. For the simplicity of presentation, we assume that the Reduce privilege refers to a single reduction function being applied (such as addition). This representation is explicitly parallel as tasks are annotated with their launch domain and partitions of distributed data structures. However, the representation is scale-free as the size of the representation is independent of the degree of parallelism — only the symbolic size of the launch domain increases.
@@ 4.3. Proof of Correctness
Given a window of tasks $[T_{1},\ldots,T_{n}]$, our task fusion algorithm identifies a prefix $[T_{1},\ldots,T_{f}]$ and produces a fused task $F$ such that
  1. (1)
    $[T_{1},\ldots,T_{f}]$ are fusible, and
  2. (2)
    $F$ preserves all dependencies of the task sequence $[T_{1},\ldots,T_{f}]$.
We provide a proof sketch for each component of the theorem. To prove that $[T_{1},\ldots,T_{f}]$ are fusible, we must show that for each pair of tasks $T_{i},T_{j},i<j$ in $[T_{1},\ldots,T_{f}]$, $\forall\leavevmode\nobreak\ p,\leavevmode\nobreak\ \mathcal{D}(T_{i},T_{j})[p]\subseteq\{p\}$. The launch-domain-equivalence constraint ensures that the dependence map is between points of the same dimensionality. For the sake of obtaining a contradiction, suppose $\exists\leavevmode\nobreak\ p,p^{\prime}$ such that $p\neq p^{\prime}$ and $\textsf{depends}(T_{i}^{p},T_{j}^{p^{\prime}})$. Then one of the three dependencies in Definition 1 must exist. Suppose that the condition for true-dep is satisfied, meaning that $\exists S,P,P^{\prime}$ such that $S[P,p]\cap S[P^{\prime},p^{\prime}]$ and $\textsf{W}(T_{i},(S,P))$ and one of $\textsf{R}(T_{j},(S,P^{\prime}))$, $\textsf{W}(T_{j},(S,P^{\prime}))$ or $\textsf{Rd}(T_{j},(S,P^{\prime})$ is true. $\textsf{R}(T_{j},(S,P^{\prime}))$ or $\textsf{W}(T_{j},(S,P^{\prime}))$ are contradictions, as the true-dependence constraint would disallow fusion. $\textsf{Rd}(T_{j},(S,P^{\prime}))$ is a contradiction due to the reduction constraint. Similar logic can be applied to other dependence cases. Here, we show that our algorithm is sound by identifying cases where fusion is possible—we do not claim completeness by proving the converse.
We have shown that all dependencies between index tasks are at most point-wise, so any $T_{j}^{p}$ can only depend on $T_{i}^{p}$, where $i<j$. Since the fused task body is the composition of each task in $[T_{1},\ldots,T_{f}]$ in program order, all dependencies in $[T_{1},\ldots,T_{f}]$ are preserved.
@@ 5.1. Temporary Store Elimination
(a) cuPyNumeric code fragment.
@@
To introduce when a store is temporary, consider the cuPyNumeric program in Figure 8(a) and the resulting task stream in Figure 8(b). This example introduces some new operations, specifically flush_window, which sends all pending tasks through Diffuse to the underlying runtime system, and the Python del operator, which drops references. The program creates the stores x, y, z, w, and v. Consider the program state after line 10: the tasks that initialize x and y have executed, as the first flush_window call sent those tasks to Diffuse. We note that there are no pending tasks outside the window, and future tasks are ones the application may launch once the call to flush_window returns. The fusion algorithm determines that the tasks issued by lines 46 can be fused, while the final norm must be excluded. First, v is not temporary because the application holds a reference to it, meaning that it could launch a task that reads v after the call to flush_window(). Next, while the application has deleted its reference to w, the norm task reads a piece of w and is still pending after the fused task, and thus must observe any effects performed on w, meaning that w is not temporary. The stores x and y are only read by the fused task, and thus are not temporary. Only z is temporary because it is produced entirely within the fused task and is not visible to the application or pending tasks. We formalize this intuition as constraints that must be satisfied for a store to be temporary.
@@
The function $\textsf{covers}(S,P)$ is true when the partition $P$ contains all points in the store $S$. The first two constraints check that the store’s contents are entirely created within the fused task and not used by any other existing task; these conditions are checked through a forwards dataflow analysis of the task stream. The third constraint ensures that the application can no longer view any effects on a store, checked through a split reference counting scheme in the implementation of Diffuse’s IR. The split reference counting scheme separates references held by the application from references held by Diffuse’s runtime. Temporary stores are demoted from a distributed allocation into a task-local allocation, as described in Section 6.
@@ 6. Kernel Fusion
The final component of Diffuse is a compilation stack to optimize fused tasks. A high-level program representation is required to both perform optimizations like loop fusion and to lower to different backends like GPUs and multi-threaded CPUs. We leverage the MLIR compiler stack, which is extensible and is pre-packaged with many common compiler analyses. We first provide background on MLIR, and then describe the code generation process and optimizations performed within Diffuse. We then discuss how Diffuse’s architecture enables the separation of reasoning about distributed programs from the optimization of nested loops.
@@ 6.2. Generator Functions
To describe Diffuse’s compiler, we walk through the stages that a fused task traverses. cuPyNumeric and Legate Sparse developers implement tasks by defining variants that target CPUs or GPUs. To use Diffuse, developers register a generator function with Diffuse that returns an MLIR fragment describing the task’s computation. We found the integration effort of adding these generator functions to be modest, requiring 50–100 lines of C++ code per operation. We emphasize that only library developers, not end users, must develop MLIR kernels for tasks. Additionally, the integration effort was incremental—as more tasks were implemented with MLIR generators, Diffuse could exploit more fusion. An example generated MLIR fragment by cuPyNumeric for an element-wise addition operation is shown in Figure 12(a).
@@ 6.3. Compilation Pipeline
When Diffuse identifies that a sequence of tasks may be fused, it invokes each task generator and constructs an MLIR module containing the body of each task in the original program order. Figure 12(b) shows a fused task for the cuPyNumeric computation c = a + b; e = c + d, where all variables represent distributed vectors. This program originally has two index tasks (one for each add operation) which are fused into a single index task where the original task bodies (the MLIR in Figure 12(a)) appear sequentially in the fused task. Before optimization of the task body, Diffuse first promotes distributed data into task-local allocations, resulting in Figure 12(c).